OpenStudy (anonymous):

The perimeter of a rectangle with length, l and width, w is given by P = 2w + 2l. If the perimeter of a certain rectangle is 78 centimeters and the width is 12 centimeters, what is the length? Please show all of your work.

OpenStudy (callisto):

P = 2w + 2l Put P=78 and w=12 into the equation 78 = 2 (12) + 2l 78 = 24 + 2l 2l = 54 solve l yourself, only 1 step is left

OpenStudy (anonymous):

how

OpenStudy (anonymous):

nevermind i got it

OpenStudy (callisto):

..... 2l =54 divide the both sides by 2 and you'll get the answer